A weak acid has a dissociation constant Ka = 2.5x10-2. Calculate the percentage dissociation for a 0.0750 m solution of this acid assuming a) ideal conditions and b) non-ideal conditions
Answer #1

dissociation of weak acid is as follows   

HA + H2O ---> H30+ +A-

I 0.075 m 0 Initial concentration

C -x +x +x change in concentration

E 0.075-x x +x Equilibrium concentration


Ka = [H3O+][A-]/[HA]

we will take equilibrium concentration, and ka =2.5*10^-2,
and plug it in above equation

2.5 *10^-2= x^2/0.075-x
lets ignore x in denominator as it is too small

so the equation become
2.5*10^-2 *0.075 =x^2
x^2 = 0.1875*10^-2
x=0.43*10^-1= 0.043

percentage dissociation = [0.043/ 0.075]*100=57.3 %

THANK YOU!!